# Quantum density estimation with density matrices: Application to quantum anomaly detection

### 22 Dec 2023 v4

@@ Appendix B Gradients and complexity of the classical and quantum QAFF
In this section, we illustrate how to find the gradients of the loss function $\mathcal{L}_{\{\mathbf{w}_{k}\}}$ that depends on the operation $\hat{k}_{\{\mathbf{w}_{k}\}}(\bm{x}_{l},\bm{x}_{m})=\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}$ for the QAFF, see Eqs. 9 and 11, in both classical and quantum computers.
@@
$\nabla_{\mathbf{w}_{j}}\mathcal{L}_{\{\mathbf{w}_{k}\}}=\frac{1}{N^{2}}\sum_{l=1}^{N}\sum_{m=1}^{N}\Big{[}-2\big{(}k(\bm{x}_{l},\bm{x}_{m})-\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}\big{)}\nabla_{\mathbf{w}_{j}}\big{(}\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}\big{)}\Big{]},$ (27)
where $\nabla_{\mathbf{w}_{j}}\big{(}\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}\big{)}$ may be written explicitly as,
@@
$\frac{1}{d}(i\sqrt{\gamma}(\bm{x}_{l}-\bm{x}_{m}))\Big{[}e^{i\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})}\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}-e^{-i\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})}\innerproduct{\psi(\bm{x}_{m})}{\psi(\bm{x}_{m})}{\psi(\bm{x}_{l})}\Big{]}.$ (28)
@@
To build this operation in a quantum computer we may use parameter-shift rule [60], by shifting the parameter $(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m}))$ of |$\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}$ to $(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})+\pi/2)$ and $(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})-\pi/2)$, since |$\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}$ might be written as $\bra{x}U^{\dagger}(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m}))\hat{B}U(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m}))\ket{x}$, for some state |$\ket{x}$ and some operator $\hat{B}$ on $n$ qubits that do not depend on $\mathbf{w}_{j}$. Hence from parameter-shift rule the gradient can also be computed by,
$\nabla_{\mathbf{w}_{j}}\big{(}\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}\big{)}=\frac{1}{2}\sqrt{\gamma}(\bm{x}_{l}-\bm{x}_{m})\Big{[}\bra{x}U^{\dagger}(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})+\pi/2)\hat{B}U(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})+\pi/2)\ket{x}$
@@
$\frac{1}{2}\sqrt{\gamma}(\bm{x}_{l}-\bm{x}_{m})\Big{[}\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}_{\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})\rightarrow\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})+\pi/2}-\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}_{\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})\rightarrow\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})-\pi/2}\Big{]},$
@@
Knowing the gradients of the loss function is sufficient to optimize the weights of the QAFF hence the total quantum complexity of QAFF lies on estimating these gradients, see Eq. 27. In Appendix A, we showed that to estimate |$\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}$ it is required $O(TDd+TdR)$ where $D$ is the size of the classical data, $d$ is the number of Fourier features, $T$ is the number of data pairs $\{(\bm{x}_{l},\bm{x}_{m})\}$, and $R$ is the number of |$\ket{0}_{n}$ measurements of the quantum kernel estimation circuit. The estimation of the gradients $\nabla_{\mathbf{w}_{j}}(\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2})$ accounts for an additional quantum complexity of $O(Td^{2}R)$ considering that for each of the $d$ Fourier vector weights $\{\mathbf{w}_{j}\}$ (or $\{\bm{\theta}_{j}\}$) and each data pair $(\bm{x}_{l},\bm{x}_{m})$ we need to find its gradient by estimating the projection |$\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}$ with shifted parameter $\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})$ with complexity $O(dR)$, because we only recalculate the angles $\{\sqrt{\gamma}\mathbf{w}_{j}\cdot\bm{x}\}$ when doing the gradient estimation of a different data pair. Therefore, the overall quantum complexity to train the QAFF corresponds to $O(TDd+Td^{2}R)$. Furthermore, the complexity of extracting the quantum Fourier features of a data set with $N$ samples is $O(NDd+Nd\log{d}R)$ ), because for each data sample, it is required $O(Dd)$ to find the Fourier angles of the QAFF, and $O(d\log{d}R)$ ) to prepare the $d$ Fourier components and to measure $R$ times the $n=\log{d}$ ) bit strings that form the basis.
@@ Appendix C SDM ansatz, density matrix preparation from eigenbasis
then, a cascade of $m$ CNOT gates between the first $m$ qubits and the last $m$ qubits are applied to the circuit, as shown in Fig. 11. The $j^{\text{th}}$ CNOT operates with control the $j^{\text{th}}$ qubit and target the $(n+j)^{\text{th}}$ qubit for $j\in\{0,\ldots m-1\}$. The resulting quantum state of this operation is,
@@
Appendix D Additional experiments on hybrid 2-D density estimation with quantum variational QAFF
@@
Although the hybrid density estimation method Q-DEMDE uses classical computers to train the quantum adaptive Fourier features, in this section, we include additional experiments to illustrate the feasibility of using a quantum simulator to learn the quantum adaptive Fourier features for density estimation via the quantum variational algorithm described in Sect. IV.2.2.
@@
Figure 12: Two-dimensional quantum-classical density estimation with the proposed Q-DEMDE algorithm with 16 Fourier components for both QRFF and QAFF. In contrast to the experiments of Sect. V.2, the QAFF were trained using a Pennylane noiseless quantum simulator. From top to bottom, the two-dimensional data sets correspond to Binomial, Potential 1, Potential 2, Arc, and Star Eight.
D.1 Experimental set-up
@@
We present the experimental results for 2-dimensional density estimation using the Q-DEMDE algorithm with the QAFF learned in a Pennylane noiseless quantum simulator. Similar to the experiments of Sect. V.2, we tested the algorithm on five 2-dimensional density estimation data sets: Binomial, Potential 1, Potential 2, Arc, and Star, shown from top to bottom in Fig. 12. The training points corresponded to 10000 points sampled from these distributions, and the test samples, whose probability density we wanted to estimate, were the result of dividing the space of each 2D data set into a grid of $120\times 120$ points.
@@
We used the quantum variational strategy described in Sect. IV.2.2 and Appendix A to train the quantum adaptive Fourier features in a noiseless quantum simulator of the Pennylane python library [67]. In total, we trained $16$ adaptive Fourier features, requiring a quantum circuit of $4$ qubits and $15\times 2=30$ variational parameters, corresponding to $15$ Fourier vector weights $\{\bm{\theta}_{j}\}\in\mathbb{R}^{2}$, using the QAFF mapping described in the Appendix A. For each 2D density estimation data set, we initialized the adaptive Fourier weights from $U[0,1]$, and used these 2-dimensional data sets as the kernel training data sets, see Sect. IV.2.1. After learning the optimal Fourier weights, $\{\bm{\theta}_{j}\}$, we prepared the QAFF mappings of the training and test density estimation data samples by making predictions from the QAFF quantum circuit. Once the QAFF were prepared, we followed the steps of the main Q-DEMDE method, i.e., we constructed the $4\times 4$ training density matrix, computed its spectral decomposition with $16$ eigenvalues, and estimated the density estimation of the test samples using the 8-qubit Q-DMKDE quantum algorithm with the Pennylane quantum simulator.
@@
We evaluated the hybrid Q-DEMDE with quantum variational QAFF by comparing the obtained probability density values with the probability densities obtained with KDE [15, 14], using the density estimation metrics KL-Div, MAE, and Spearman correlation. As a baseline, we also performed experiments using the Q-DEMDE method with 16 QRFF.
@@ D.2 Results and discussion
We present the results of the Q-DEMDE density estimation method with quantum variational QAFF in Fig. 12 and Table 10. The experimental results show that it is possible to perform quantum-classical density estimation by learning the QAFF feature map in a quantum simulator, for instance, the QDE method achieved satisfactory density estimation results, especially on the Potential 1 and the Potential 2 data sets; it is worth highlighting that the adaptive Fourier weights were randomly initialized from $U[0,1]$, thus the method was able to capture the optimal Fourier parameters. However, the results also show better performance of the Q-DEMDE method with QRFF on the Binomial, Arc and Star data sets, indicating that the results of Q-DEMDE method could have been improved by increasing the number of training iterations and using the Gaussian kernel training data set, see Sect. IV.2.1, which is to be explored in future work.
@@ 16 QRFF 16 QAFF
Table 10: 2D quantum-classical density estimation results with Q-DEMDE algorithm with 16 Fourier components for both QRFF and QAFF. The QAFF were trained using a Pennylane noiseless quantum simulator. The best results are in bold.
Appendix E Two Dimensional Data sets

### 18 Mar 2024 v5

@@ Appendix B Gradients and complexity of the classical and quantum QAFF
In this section, we illustrate how to find the gradients of the loss function $\mathcal{L}_{\{\mathbf{w}_{k}\}}$ that depends on the operation $\hat{k}_{\{\mathbf{w}_{k}\}}(\bm{x}_{l},\bm{x}_{m})=\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}$ for the QAFF, see Eqs. 9 and 11, in both classical and quantum computers.
@@
$\nabla_{\mathbf{w}_{j}}\mathcal{L}_{\{\mathbf{w}_{k}\}}=\frac{1}{N^{2}}\sum_{l=1}^{N}\sum_{m=1}^{N}\Big{[}-2\big{(}k(\bm{x}_{l},\bm{x}_{m})-\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}\big{)}\nabla_{\mathbf{w}_{j}}\big{(}\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}\big{)}\Big{]},$ (27)
where $\nabla_{\mathbf{w}_{j}}\big{(}\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}\big{)}$ may be written explicitly as,
@@
$\frac{1}{d}(i\sqrt{\gamma}(\bm{x}_{l}-\bm{x}_{m}))\Big{[}e^{i\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})}\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle-e^{-i\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})}\langle\psi(\bm{x}_{m})|\psi(\bm{x}_{l})\rangle\Big{]}.$ (28)
@@
To build this operation in a quantum computer we may use parameter-shift rule [61], by shifting the parameter $(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m}))$ of |$\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}$ to $(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})+\pi/2)$ and $(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})-\pi/2)$, since |$\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}$ might be written as $\bra{x}U^{\dagger}(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m}))\hat{B}U(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m}))\ket{x}$, for some state |$\ket{x}$ and some operator $\hat{B}$ on $n$ qubits that do not depend on $\mathbf{w}_{j}$. Hence from parameter-shift rule the gradient can also be computed by,
$\nabla_{\mathbf{w}_{j}}\big{(}\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}\big{)}=\frac{1}{2}\sqrt{\gamma}(\bm{x}_{l}-\bm{x}_{m})\Big{[}\bra{x}U^{\dagger}(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})+\pi/2)\hat{B}U(\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})+\pi/2)\ket{x}$
@@
$\frac{1}{2}\sqrt{\gamma}(\bm{x}_{l}-\bm{x}_{m})\Big{[}\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}_{\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})\rightarrow\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})+\pi/2}-\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}_{\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})\rightarrow\sqrt{\gamma}\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})-\pi/2}\Big{]},$
@@
Knowing the gradients of the loss function is sufficient to optimize the weights of the QAFF hence the total quantum complexity of QAFF lies on estimating these gradients, see Eq. 27. In Appendix A, we showed that to estimate |$\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}$ it is required $O(TDd+TdR)$ where $D$ is the size of the classical data, $d$ is the number of Fourier features, $T$ is the number of data pairs $\{(\bm{x}_{l},\bm{x}_{m})\}$, and $R$ is the number of |$\ket{0}_{n}$ measurements of the quantum kernel estimation circuit. The estimation of the gradients $\nabla_{\mathbf{w}_{j}}(\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2})$ accounts for an additional quantum complexity of $O(Td^{2}R)$ considering that for each of the $d$ Fourier vector weights $\{\mathbf{w}_{j}\}$ (or $\{\bm{\theta}_{j}\}$) and each data pair $(\bm{x}_{l},\bm{x}_{m})$ we need to find its gradient by estimating the projection |$\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}$ with shifted parameter $\mathbf{w}_{j}\cdot(\bm{x}_{l}-\bm{x}_{m})$ with complexity $O(dR)$, because we only recalculate the angles $\{\sqrt{\gamma}\mathbf{w}_{j}\cdot\bm{x}\}$ when doing the gradient estimation of a different data pair. Therefore, the overall quantum complexity to train the QAFF corresponds to $O(TDd+Td^{2}R)$. Furthermore, the complexity of extracting the quantum Fourier features of a data set with $N$ samples is $O(NDd+Nd\log{d}R)$ ), because for each data sample, it is required $O(Dd)$ to find the Fourier angles of the QAFF, and $O(d\log{d}R)$ ) to prepare the $d$ Fourier components and to measure $R$ times the $n=\log{d}$ ) bit strings that form the basis.
@@ Appendix C SDM ansatz, density matrix preparation from eigenbasis
then, a cascade of $m$ CNOT gates between the first $m$ qubits and the last $m$ qubits are applied to the circuit, as shown in Fig. 11. The $j^{\text{th}}$ CNOT controls the $j$th qubit and targets the $(n+j)$th qubit for $j\in\{0,\ldots m-1\}$. The resulting quantum state of this operation is,
@@
Appendix D Additional simulations for hybrid 2-D density estimation with quantum variational QAFF
@@
Although the hybrid density estimation method Q-DEMDE uses classical computers to train the quantum adaptive Fourier features, in this section, we demonstrate the feasibility of using a quantum simulator to learn the quantum adaptive Fourier features for density estimation via the quantum variational algorithm described in Sect. IV.2.2.
@@
Figure 12: Two-dimensional quantum-classical density estimation with the proposed Q-DEMDE algorithm with 16 Fourier components for both QRFF and QAFF. In contrast to the results of Sect. V.2, which use classical hardware to train the QAFF, here, we present the results of using a Pennylane noiseless quantum simulator to optimize the Fourier features. From top to bottom, the two-dimensional data sets correspond to Binomial, Potential 1, Potential 2, Arc, and Star Eight.
D.1 Setup
@@
We present the results for 2-dimensional density estimation using the Q-DEMDE algorithm with the QAFF learned in a Pennylane noiseless quantum simulator. Similar to the simulations of Sect. V.2, we tested the algorithm on five 2-dimensional density estimation data sets: Binomial, Potential 1, Potential 2, Arc, and Star, shown from top to bottom in Fig. 12. The training points corresponded to 10000 points sampled from these distributions, and the test samples, whose probability density we wanted to estimate, were the result of dividing the space of each 2D data set into a grid of $120\times 120$ points.
@@
We used the quantum variational strategy described in Sect. IV.2.2 and Appendix A to train the quantum adaptive Fourier features in a noiseless quantum simulator of the Pennylane python library [68]. In total, we trained $16$ adaptive Fourier features, requiring a quantum circuit of $4$ qubits and $15\times 2=30$ variational parameters, corresponding to $15$ Fourier vector weights $\{\bm{\theta}_{j}\}\in\mathbb{R}^{2}$, using the QAFF mapping described in the Appendix A. For each 2D density estimation data set, we initialized the adaptive Fourier weights from $U[0,1]$, and used these 2-dimensional data sets as the kernel training data sets, see Sect. IV.2.1. After learning the optimal Fourier weights, $\{\bm{\theta}_{j}\}$, we prepared the QAFF mappings of the training and test density estimation data samples by making predictions from the QAFF quantum circuit. Once the QAFF were prepared, we followed the steps of the main Q-DEMDE method, i.e., we constructed the $16\times 16$ training density matrix, computed its spectral decomposition with $16$ eigenvalues, and estimated the density estimation of the test samples using the 8-qubit Q-DMKDE quantum algorithm with the Pennylane quantum simulator.
@@
We evaluated the hybrid Q-DEMDE with quantum variational QAFF by comparing the obtained probability density values with the probability densities obtained with KDE [15, 14], using the density estimation metrics KL-Div, MAE, and Spearman correlation. As a baseline, we also performed hybrid density estimation using the Q-DEMDE method with 16 QRFF.
@@ D.2 Results and discussion
We present the results of the Q-DEMDE density estimation method with quantum variational QAFF in Fig. 12 and Table 10. The results of the simulations show that it is possible to perform quantum-classical density estimation by learning the QAFF feature map in a quantum simulator, for instance, the QDE method achieved satisfactory results for density estimation, especially on the Potential 2 data set; it is worth highlighting that the adaptive Fourier weights were randomly initialized from $U[0,1]$, thus the method was able to capture the optimal Fourier parameters. However, the results also show better performance of the Q-DEMDE method with QRFF on the Binomial, Potential 1, Arc and Star data sets, indicating that the results of Q-DEMDE method could have been improved by increasing the number of training iterations and using the Gaussian kernel training data set, see Sect. IV.2.1, which is to be explored in future work.
@@ 16 QRFF 16 QAFF
Table 10: 2D quantum-classical density estimation results with Q-DEMDE algorithm with 16 Fourier components for both QRFF and QAFF. The QAFF were trained using a Pennylane noiseless quantum simulator. The best results are shown in bold.
Appendix E Description of the IBM-Oslo quantum device
All the demonstrations related to applications of the methods presented in this paper on real quantum systems were performed on the IBM-Oslo quantum computer, one of the quantum processors made publicly available through the IBM Quantum Platform. The layout of the quantum processor is shown in Fig. 13, and other relevant features of the processor are listed in Table 11.
Figure 13: Topological map of the IBM-Oslo quantum processor.
CLOPS (Circuit layer operations per second) 2.6K
Median fundamental transition frequency 5.046 GHz
Median anharmonicity -0.3429 GHz
Median qubit lifetime $T_{1}$ 128 $\mu$s
Median coherence time $T_{2}$ 79.89 $\mu$s
Single qubit gate error $1.65\times 10^{-4}$ to $6.698\times 10^{-4}$
CNOT gate error $6.47\times 10^{-3}$ to $2.067\times 10^{-2}$
Average readout error 0.0216
Table 11: Main features of the IBM-Oslo quantum processor. The IBM-Oslo parameters were taken from Ref. [77].
Appendix F Two-dimensional data sets